A particle of mass m moves with momentum of magnitude p. (a) Show that the kinetic energy of the particle is K = p^2/ 2m. (b) An object has a kinetic energy of 275 J and momentum of magnitude 25.0 kgm/s. Find the
and mass of the object A
1
Expert's answer
2020-11-30T15:23:40-0500

Answer

a) kinetic energy is 

"KE=\frac{mv^2}{2}"

Momentum is given by

"P= mv\\v=\frac{P}{m}"

Putting in kinetic energy become as

"KE=\frac{mP^2}{2m^2}=\frac{P^2}{2m}" Hence proved

b) mass is given by

"m=\frac{P^2}{2K}=\frac{(25)^2}{2\times275}=1.136Kg"